OpenStudy (anonymous):

1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = [n(6n^2 - 3n - 1)]/2

OpenStudy (zzr0ck3r):

induction?

OpenStudy (anonymous):

Yeah. :/

OpenStudy (zzr0ck3r):

sec

OpenStudy (anonymous):

Okay! :)

OpenStudy (zzr0ck3r):

would take way to long to type it. ask if you have questions. note how easy it is if you use summation notation.

OpenStudy (anonymous):

Thank you so much. I've been working on this problem all day! I think that I actually understand induction based on how you wrote it out! Thank you so much again!

OpenStudy (anonymous):

u LARS ?

OpenStudy (anonymous):

Who the hell are u?

OpenStudy (anonymous):

You are such a loser ** get a life *

OpenStudy (anonymous):

Bye!

OpenStudy (anonymous):

@_Unknown_17 Excuse me... Who are you?

OpenStudy (zzr0ck3r):

np @Mrbonez321 just think of dominoes. this method will work on all the summation ones, and will help you get a better understanding of induction, i agree.

OpenStudy (anonymous):

@zzr0ck3r Thank you for your help. I just have a quick question. On the foil, you got 6n^3 + 15n^2 + 11n + 2... But i dont understand where you got the 15n^2 from...

OpenStudy (anonymous):

Ohhh i don't wanna talk! bye !

OpenStudy (zzr0ck3r):

dont worry about the douchebags, there are plenty on the site.

OpenStudy (zzr0ck3r):

on the rhs or the lhs?

OpenStudy (anonymous):

LHS

OpenStudy (zzr0ck3r):

are you talking about after this step 9n^2+6n+1+(n(6n^2-3n-1)/2)

OpenStudy (anonymous):

Yes.

OpenStudy (zzr0ck3r):

(9n^2+6n+1)+(n(6n^2-3n-1)/2) we need a common denom of 2, so we multiply 9n^2+6n+1 (top/bottom) by 2 (18n^2+12n+2)/2 now we have (18n^2+12n+2)/2 +(n(6n^2-3n-1)/2) = [(18n^2+12n+2)+(n(6n^2-3n-1)]/2 = [18n^2+12n+2+6n^3-3n^2-n]/2 = [6n^3 + 18n^2-3n^2+12n-11n+2]/2 = [6n^3+15n+11n+2]/2

OpenStudy (anonymous):

Oh!!!! I forgot to multiply by two on the 9n^2 +6n+1 part!!! Thank you so much!!!

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

a+(b/c) = (ac+b)/c